Every single day I will be delivered a random number of fanmail in the mail. I will sanitize each one with my hand sanitizer for safety. I have enough sanitizer to sanitize 600 fanmail. Each and every day, I get at least 2 fanmail, but not more than 12, and I always receive an even number of fanmail. Among these possible numbers of fanmail I may possibily receive on any given day, all are equally likely to happen. What is the approximate probability that my stockpile of hand sanitizer will last atleast 2 months (For our purposes there are 30 days in every month, and assume you get mail every single day.) You can use a basic calculator.

If you can explain you reasoning. Show all work!!!! Don't leave steps out please.
Answer #1

Solution:-

X = Random number fanmail in the mail

p(X = 2) = p(X = 4) + p(X = 6) + p(X = 8) + p(X = 10) + p(X = 12)

p(X = 2) =

  

  

  

   ( n = 90)

p(sanities last at least 2 months)

= p(at tea most 600 fanmail in 90 days)

= 0.1772